What will be the radius of curvature of the path of a 3.0 KeV proton in a

perpendicular magnetic field of magnitude 0.8 T ?


1
Expert's answer
2021-05-30T13:30:35-0400

Larmor radius is given by a formula:

"R = \frac {mv}{qB}"

The energy of electron:

"E = \frac{mc^2}{1-(\frac v c)^2}"

In this case we can express a speed of electron:

"v=c\sqrt{\frac{E-mc^2}{E}} = 534976 \frac m s"

"\frac m q = 10^{-3}"

"R = 668 m"